# Bandits with Preference Feedback: A Stackelberg Game Perspective

### v2

@@ 1 Introduction
In standard bandit optimization, a learner repeatedly interacts with an unknown environment that gives numerical feedback on the chosen actions according to a utility function $f$. However, in applications such as fine-tuning large language models, drug testing, or search engine optimization, the quantitative value of design choices or test outcomes are either not directly observable, or are known to be inaccurate, or systematically biased, e.g., if they are provided by human feedback (Casper et al., 2023). A solution is to optimize for the target based on comparative feedback provided for a pair of queries, which is proven to be more robust to certain biases and uncertainties in the queries (Ji et al., 2023).
Bandits with preference feedback, or dueling bandits, address this problem and propose strategies for choosing query/action pairs that yield a high utility over the horizon of interactions. At the core of such algorithms is uncertainty quantification and inference for $f$ in regions of interest, which is closely tied to exploration and exploitation dilemma over a course of queries. Observing only comparative feedback poses an additional challenge, as we now need to balance this trade-off jointly over two actions. This challenge is further exacerbated when optimizing over vast or infinite action domains. As a remedy, prior work often grounds one of the actions by choosing it either randomly or greedily, and tries to balance exploration-exploitation for the second action as a reaction to the first (Ailon et al., 2014; Zoghi et al., 2014a; Kirschner and Krause, 2021; Mehta et al., 2023b). This approach works well for simple utility functions over low-dimensional domains, however does not scale to more complex problems.
@@ 2 Related Work
Learning with indirect feedback was first studied in the supervised preference learning setting (Aiolli and Sperduti, 2004; Chu and Ghahramani, 2005). Subsequently, online and sequential settings were considered, motivated by applications in which the feedback is provided in an online manner, e.g., by a human (Yue et al., 2012; Yue and Joachims, 2009; Houlsby et al., 2011). Bengs et al. (2021) surveys this field comprehensively; here we include a brief background.
@@
Referred to as dueling bandits, a rich body of work considers (finite) multi-armed domains and learns a preference matrix specifying the relation among the arms. Such work often relies on efficient sorting or tournament systems based on the frequency of wins for each action (Jamieson and Nowak, 2011; Zoghi et al., 2014b; Komiyama et al., 2015; Wu and Liu, 2016; Falahatgar et al., 2017). Rather than jointly selecting the arms, such strategies often simplify the problem by selecting one at random (Zoghi et al., 2014a; Zimmert and Seldin, 2018), greedily (Chen and Frazier, 2017), or from the set of previously selected arms (Ailon et al., 2014). In contrast, we jointly optimize both actions by choosing them as the equilibrium of a two-player zero-sum Stackelberg game, enabling a more efficient exploration/exploitation trade-off.
@@
The multi-armed dueling setting, which is reducible to multi-armed bandits (Ailon et al., 2014), naturally fails to scale to infinite compact domains, since regularity among “similar” arms is not exploited. To go beyond finite domains, utility-based dueling bandits consider an unknown latent function that captures the underlying preference, instead of relying on a preference matrix. The preference feedback is then modeled as the difference in the utility of two chosen actions passed through a link function. Early work is limited to convex domains and imposes strong regularity assumptions (Yue and Joachims, 2009; Kumagai, 2017). These assumptions are then relaxed to general compact domains if the utility function is linear (Dudík et al., 2015; Saha, 2021; Saha and Krishnamurthy, 2022). Constructing valid confidence sets from comparative feedback is a challenging task. However, it is strongly related to uncertainty quantification with direct logistic feedback, which is extensively analyzed by the literature on logistic and generalized linear bandits (Filippi et al., 2010; Faury et al., 2020; Foster and Krishnamurthy, 2018; Beygelzimer et al., 2019; Faury et al., 2022; Lee et al., 2024).
@@
Preference-based bandit optimization with linear utility functions is well understood and is even extended to reinforcement learning with preference feedback on trajectories (Saha et al., 2023; Zhan et al., 2023; Zhu et al., 2023; Ji et al., 2023; Munos et al., 2023). However, such approaches have limited practical interest, since they cannot capture real-world problems with complex nonlinear utility functions. Alternatively, Reproducing Kernel Hilbert Spaces (RKHS) provide a rich model class for the utility, e.g., if the chosen kernel is universal. Many have proposed heuristic algorithms for bandits and Bayesian optimization in kernelized settings, albeit without providing theoretical guarantees Brochu et al. (2010); González et al. (2017); Sui et al. (2017); Tucker et al. (2020); Mikkola et al. (2020); Takeno et al. (2023).
@@
There have been attempts to prove convergence of kernelized algorithms for preference-based bandits (Xu et al., 2020; Kirschner and Krause, 2021; Mehta et al., 2023b, a). Such works employ a regression likelihood model which requires them to assume that both the utility and the probability of preference, as a function of actions, lie in an RKHS. In doing so, they use a regression model for solving a problem that is inherently a classification. While the model is valid, it does not result in a sample-efficient algorithm. In contrast, we use a kernelized logistic negative log-likelihood loss to infer the utility function, and provide confidence sets for its minimizer. In a concurrent work, Xu et al. (2024) also consider the kernelized logistic likelihood model and propose a variant of the MultiSBM algorithm (Ailon et al., 2014) which uses likelihood ratio confidence sets. The theoretical approach and resulting algorithm bear significant differences, and the regret guarantee has a strictly worse dependency on the time horizon $T$, by a factor of $T^{1/4}$. This is discussed in more detail in Section 5.
@@ 3 Problem Setting
which is the average sub-optimality gap between the chosen pair and a globally preferred action ${\bm{x}}^{\star}$. To better understand this notion of regret, consider the scenario where actions ${\bm{x}}_{t}$ and ${\bm{x}}_{t}^{\prime}$ are both optimal. Then the probabilities are equal to $0.5$ and the dueling regret will not grow further, since the regret incurred at step $t$ is zero. This formulation of $R^{\mathrm{D}}(T)$ is commonly used in the literature of dueling Bandits and RL with preference feedback (Urvoy et al., 2013; Saha et al., 2023; Zhu et al., 2023) and is adapted from Yue et al. (2012). Our goal is to design an algorithm that satisfies a sublinear dueling regret, where $R^{\mathrm{D}}(T)/T\rightarrow 0$ as $T\rightarrow\infty$. This implies that given enough evidence, the algorithm will converge to a globally preferred action. To this end, we take a utility-based approach and consider an unknown utility function ${f}:{\mathcal{X}}\rightarrow{\mathbb{R}}$, which reflects the preference via
@@
where $s:{\mathbb{R}}\rightarrow[0,1]$ is the sigmoid function [^1] May be generalized to differentiable monotonically increasing functions satisfying $s(x)+s(-x)=1$., i.e. $s(a)=(1+e^{-a})^{-1}$. Referred to as the Bradley-Terry model (Bradley and Terry, 1952), this probabilistic model for binary feedback is widely used in the literature for logistic and generalized bandits (Filippi et al., 2010; Faury et al., 2020). Under the utility-based model, ${\bm{x}}^{\star}=\operatorname*{arg\,max}_{{\bm{x}}\in{\mathcal{X}}}f({\bm{x}})$ and we can draw connections to a classic bandit problem with direct feedback over the utility $f$. In particular, Saha (2021) shows that the dueling regret of (1) is equivalent up to constant factors, to the average utility regret of the two actions, that is $\sum_{t=1}^{T}{f}({\bm{x}}^{\star})-[{f}({\bm{x}}_{t})+{f}({\bm{x}}_{t}^{\prime})]/2$.
@@
While our theoretical framework targets euclidean domains, our methodology may be used on general domains of text or images, given vector embeddings obtained via unsupervised learning. Solving a bandit problem on top of embeddings from a pretrained language model is common practice in further fine-tuning of such models (e.g., Nguyen et al., 2024; Mehta et al., 2023a), and is further demonstrated in our Yelp experiment (c.f Section 6.3). Lastly, we highlight that our results may be smoothly extended to contextual bandits with stochastic context, by simply modifying the signature of the kernel function to $k^{\prime}({\bm{x}},{\bm{x}}^{\prime},{\bm{z}}):{\mathcal{X}}\times{\mathcal{X}}\times{\mathcal{Z}}\rightarrow{\mathbb{R}}$, where ${\bm{z}}\in{\mathcal{Z}}$ denotes the context. This setting accommodates applications in active learning for fine-tuning of large language models, where the context is the prompt and the two actions are two alternative responses.
@@ 4 Kernelized Confidence Sequences with Direct Logistic Feedback
As a warm-up, we consider a hypothetical scenario where ${\bm{x}}^{\prime}_{t}={\bm{x}}_{\mathrm{null}}$ for all $t\geq 1$ such that ${f}({\bm{x}}_{\mathrm{null}})=0$. Therefore at every step, we suggest an action ${\bm{x}}_{t}$ and receive a noisy binary feedback $y_{t}$, which is equal to one with probability $s({f}({\bm{x}}_{t}))$. This example reduces our problem to logistic bandits which has been rigorously analyzed for linear rewards (Filippi et al., 2010; Faury et al., 2020). We extend prior work to the non-parametric setting by proposing a tractable loss function for estimating the utility function, a.k.a. reward. We present novel confidence intervals that quantify the uncertainty on the logistic predictions uniformly over the action domain. In doing so, we propose confidence sequences for the kernelized logistic likelihood model that are of independent interest for developing sample-efficient solvers for online and active classification.
The feedback $y_{t}$ is a Bernoulli random variable, and its likelihood depends on the utility function as $s(f({\bm{x}}_{t}))^{y_{t}}[1-s(f({\bm{x}}_{t}))]^{1-y_{t}}$. Then given history $H_{t}$, we can estimate ${f}$ by ${f_{t}}$, the minimizer of the regularized negative log-likelihood loss
@@
Solving for ${f_{t}}$ may seem intractable at first glance since the loss is defined over functions in the large space of ${\mathcal{H}}_{k}$. However, it is common knowledge that the solution has a parametric form and may be calculated by using gradient descent. This is a direct application of the Representer Theorem (Schölkopf et al., 2001) and is detailed in Proposition 1.
@@
Given ${f_{t}}$, we may predict the expected feedback for a point ${\bm{x}}$ as $s({f_{t}}({\bm{x}}))$. Centered around this prediction, we construct confidence sets of the form $[s({f_{t}}({\bm{x}}))\pm\beta_{t}(\delta)\sigma_{t}({\bm{x}})]$, and show their uniform anytime validity. The width of the sets are characterized by $\sigma_{t}({\bm{x}})$ defined as
@@
where $\kappa=\sup_{a\leq B}1/\dot{s}(a)$, with $\dot{s}(a)=s(a)(1-s(a))$ denoting the derivative of the sigmoid function, and $K_{t}\in{\mathbb{R}}^{t\times t}$ is the kernel matrix satisfying $[K_{t}]_{i,j}=k({\bm{x}}_{i},{\bm{x}}_{j})$. Our first main result shows that for a careful choice of $\beta_{t}(\delta)$, these sets contain $s(f({\bm{x}}))$ simultaneously for all ${\bm{x}}\in{\mathcal{X}}$ and $t\geq 1$ with probability greater than $1-\delta$.
@@
Assume $f\in{\mathcal{H}}_{k}$ and $\left\lVert f\right\rVert_{k}\leq B$. Let $10<\delta<$ 1 and set
@@
Function-valued confidence sets around the kernelized ridge estimator are analyzed and used extensively to design bandit algorithms with noisy feedback on the true reward values (Valko et al., 2013; Srinivas et al., 2010; Chowdhury and Gopalan, 2017; Whitehouse et al., 2023). However, under noisy logistic feedback, this literature falls short as the proposed confidence sets are no longer valid for the kernelized logistic estimator ${f_{t}}$. One could still estimate $f$ using a kernelized ridge estimator and benefit from this line of work. However, as empirically demonstrated in Figure 1(a), this will not be a sample-efficient approach.
Proof Sketch. When minimizing the kernelized logistic loss, we do not have a closed-form solution for ${f_{t}}$, and can only formulate it using the fact that the gradient of the loss evaluated at ${f_{t}}$ is the null operator, i.e., $\nabla{\mathcal{L}}({f_{t}};H_{t}):{\mathcal{H}}\rightarrow{\mathcal{H}}=\bm{0}$. The key idea of our proof is to construct confidence intervals as ${\mathcal{H}}$-valued ellipsoids in the gradient space and show that the gradient operator evaluated at $f$ belongs to it with high probability (c.f. Lemma 8). We then translate this back into intervals around point estimates $s({f_{t}}({\bm{x}}))$ uniformly for all points ${\bm{x}}\in{\mathcal{X}}$. The complete proof is deferred to Appendix A, and builds on the results of Faury et al. (2020) and Whitehouse et al. (2023).
Logistic Bandits. Such confidence sets are an integral tool for action selection under uncertainty, and bandit algorithms often rely on them to balance exploration against exploitation. To demonstrate how Theorem 2 may be used for bandit optimization with direct logistic feedback, we consider LGP-UCB, the kernelized Logistic GP-UCB algorithm. Presented in Algorithm 2, it extends the optimistic algorithm of Faury et al. (2020) from the linear to the kernelized setting, by using the confidence bound of Theorem 2 to calculate an optimistic estimate of the reward. We proceed to show that LGP-UCB attains a sublinear logistic regret, which is commonly defined as
@@
Let $\delta\in(0,1]$ and choose the exploration coefficients $\beta_{t}(\delta)$ as described in Theorem 2 for all $t\geq 0$. Then LGP-UCB satisfies the anytime cumulative regret guarantee of

### v3

@@ 1 Introduction
In standard bandit optimization, a learner repeatedly interacts with an unknown environment that gives numerical feedback on the chosen actions according to a utility function $f$. However, in applications such as fine-tuning large language models, drug testing, or search engine optimization, the quantitative value of design choices or test outcomes are either not directly observable, or are known to be inaccurate, or systematically biased, e.g., if they are provided by human feedback (casper2023open). A solution is to optimize for the target based on comparative feedback provided for a pair of queries, which is proven to be more robust to certain biases and uncertainties in the queries (ji2023provable).
Bandits with preference feedback, or dueling bandits, address this problem and propose strategies for choosing query/action pairs that yield a high utility over the horizon of interactions. At the core of such algorithms is uncertainty quantification and inference for $f$ in regions of interest, which is closely tied to exploration and exploitation dilemma over a course of queries. Observing only comparative feedback poses an additional challenge, as we now need to balance this trade-off jointly over two actions. This challenge is further exacerbated when optimizing over vast or infinite action domains. As a remedy, prior work often grounds one of the actions by choosing it either randomly or greedily, and tries to balance exploration-exploitation for the second action as a reaction to the first (ailon2014reducing; zoghi2014relative; kirschner2021bias; mehta2023kernelized). This approach works well for simple utility functions over low-dimensional domains, however does not scale to more complex problems.
@@ 2 Related Work
Learning with indirect feedback was first studied in the supervised preference learning setting (aiolli2004learning; chu2005preference). Subsequently, online and sequential settings were considered, motivated by applications in which the feedback is provided in an online manner, e.g., by a human (yue2012k; yue2009interactively; houlsby2011bayesian). bengs2021preference surveys this field comprehensively; here we include a brief background.
@@
Referred to as dueling bandits, a rich body of work considers (finite) multi-armed domains and learns a preference matrix specifying the relation among the arms. Such work often relies on efficient sorting or tournament systems based on the frequency of wins for each action (jamieson2011active; zoghi2014relative2; komiyama2015optimal; wu2016double; falahatgar2017maximum). Rather than jointly selecting the arms, such strategies often simplify the problem by selecting one at random (zoghi2014relative; zimmert2018factored), greedily (chen2017dueling), or from the set of previously selected arms (ailon2014reducing). In contrast, we jointly optimize both actions by choosing them as the equilibrium of a two-player zero-sum Stackelberg game, enabling a more efficient exploration/exploitation trade-off.
@@
The multi-armed dueling setting, which is reducible to multi-armed bandits (ailon2014reducing), naturally fails to scale to infinite compact domains, since regularity among “similar” arms is not exploited. To go beyond finite domains, utility-based dueling bandits consider an unknown latent function that captures the underlying preference, instead of relying on a preference matrix. The preference feedback is then modeled as the difference in the utility of two chosen actions passed through a link function. Early work is limited to convex domains and imposes strong regularity assumptions (yue2009interactively; kumagai2017regret). These assumptions are then relaxed to general compact domains if the utility function is linear (dudik2015contextual; saha2021optimal; saha2022efficient). Constructing valid confidence sets from comparative feedback is a challenging task. However, it is strongly related to uncertainty quantification with direct logistic feedback, which is extensively analyzed by the literature on logistic and generalized linear bandits (filippi2010parametric; faury2020improved; foster2018contextual; beygelzimer2019bandit; faury2022jointly; lee2024improved).
@@
Preference-based bandit optimization with linear utility functions is well understood and is even extended to reinforcement learning with preference feedback on trajectories (pacchiano2021dueling; zhan2023provable; zhu2024principled; ji2023provable; munos2023nash). However, such approaches have limited practical interest, since they cannot capture real-world problems with complex nonlinear utility functions. Alternatively, Reproducing Kernel Hilbert Spaces (RKHS) provide a rich model class for the utility, e.g., if the chosen kernel is universal. Many have proposed heuristic algorithms for bandits and Bayesian optimization in kernelized settings, albeit without providing theoretical guarantees brochu2010tutorial; gonzalez2017preferential; sui2017multi; tucker2020preference; mikkola2020projective; takeno2023towards.
@@
There have been attempts to prove convergence of kernelized algorithms for preference-based bandits (xu2020zeroth; kirschner2021bias; mehta2023kernelized; mehta2023sample). Such works employ a regression likelihood model which requires them to assume that both the utility and the probability of preference, as a function of actions, lie in an RKHS. In doing so, they use a regression model for solving a problem that is inherently a classification. While the model is valid, it does not result in a sample-efficient algorithm. In contrast, we use a kernelized logistic negative log-likelihood loss to infer the utility function, and provide confidence sets for its minimizer. In a concurrent work, xu2024principled also consider the kernelized logistic likelihood model and propose a variant of the MultiSBM algorithm (ailon2014reducing) which uses likelihood ratio confidence sets. The theoretical approach and resulting algorithm bear significant differences, and the regret guarantee has a strictly worse dependency on the time horizon $T$, by a factor of $T^{1/4}$. This is discussed in more detail in Section˜5.
@@ 3 Problem Setting
which is the average sub-optimality gap between the chosen pair and a globally preferred action $\bm{x}$$^{\star}$. To better understand this notion of regret, consider the scenario where actions ${\bm{x}}_{t}$ and ${\bm{x}}_{t}^{\prime}$ are both optimal. Then the probabilities are equal to $0.5$ and the dueling regret will not grow further, since the regret incurred at step $t$ is zero. This formulation of $R^{\mathrm{D}}(T)$ is commonly used in the literature of dueling Bandits and RL with preference feedback (urvoy2013generic; pacchiano2021dueling; zhu2024principled) and is adapted from yue2012k. Our goal is to design an algorithm that satisfies a sublinear dueling regret, where $R^{\mathrm{D}}(T)/T\rightarrow 0$ as $T\rightarrow\infty$. This implies that given enough evidence, the algorithm will converge to a globally preferred action. To this end, we take a utility-based approach and consider an unknown utility function ${f}:{\mathcal{X}}\rightarrow{\mathbb{R}}$, which reflects the preference via
@@
where $s:{\mathbb{R}}\rightarrow[0,1]$ is the sigmoid function [^1] May be generalized to differentiable monotonically increasing functions satisfying s(x)+s(x)=1s(x)+s(-x)=1., i.e. $s(a)=(1+e^{-a})^{-1}$. Referred to as the Bradley-Terry model (bradley1952rank), this probabilistic model for binary feedback is widely used in the literature for logistic and generalized bandits (filippi2010parametric; faury2020improved). Under the utility-based model, ${\bm{x}}^{\star}=\operatorname*{arg\,max}_{{\bm{x}}\in{\mathcal{X}}}f({\bm{x}})$ and we can draw connections to a classic bandit problem with direct feedback over the utility $f$. In particular, saha2021optimal shows that the dueling regret of (1) is equivalent up to constant factors, to the average utility regret of the two actions, that is $\sum_{t=1}^{T}{f}({\bm{x}}^{\star})-[{f}({\bm{x}}_{t})+{f}({\bm{x}}_{t}^{\prime})]/2$.
@@
While our theoretical framework targets euclidean domains, our methodology may be used on general domains of text or images, given vector embeddings obtained via unsupervised learning. Solving a bandit problem on top of embeddings from a pretrained language model is common practice in further fine-tuning of such models (e.g., nguyen2024predicting; mehta2023sample), and is further demonstrated in our Yelp experiment (c.f Section˜6.3). Lastly, we highlight that our results may be smoothly extended to contextual bandits with stochastic context, by simply modifying the signature of the kernel function to $k^{\prime}({\bm{x}},{\bm{x}}^{\prime},{\bm{z}}):{\mathcal{X}}\times{\mathcal{X}}\times{\mathcal{Z}}\rightarrow{\mathbb{R}}$, where ${\bm{z}}\in{\mathcal{Z}}$ denotes the context. This setting accommodates applications in active learning for fine-tuning of large language models, where the context is the prompt and the two actions are two alternative responses.
@@ 4 Kernelized Confidence Sequences with Direct Logistic Feedback
As a warm-up, we consider a hypothetical scenario where ${\bm{x}}^{\prime}_{t}={\bm{x}}_{\mathrm{null}}$ for all $t\geq 1$ such that ${f}({\bm{x}}_{\mathrm{null}})=0$. Therefore at every step, we suggest an action ${\bm{x}}_{t}$ and receive a noisy binary feedback $y_{t}$, which is equal to one with probability $s({f}({\bm{x}}_{t}))$. This example reduces our problem to logistic bandits which has been rigorously analyzed for linear rewards (filippi2010parametric; faury2020improved). We extend prior work to the non-parametric setting by proposing a tractable loss function for estimating the utility function, a.k.a. reward. We present novel confidence intervals that quantify the uncertainty on the logistic predictions uniformly over the action domain. In doing so, we propose confidence sequences for the kernelized logistic likelihood model that are of independent interest for developing sample-efficient solvers for online and active classification.
The feedback $y_{t}$ is a Bernoulli random variable, and its likelihood depends on the utility function as $s(f({\bm{x}}_{t}))^{y_{t}}[1-s(f({\bm{x}}_{t}))]^{1-y_{t}}$. Then given history $H_{t}$, we can estimate f{f} by $f_{t}$}, the minimizer of the regularized negative log-likelihood loss
@@
Solving for $f_{t}$} may seem intractable at first glance since the loss is defined over functions in the large space of ${\mathcal{H}}_{k}$. However, it is common knowledge that the solution has a parametric form and may be calculated by using gradient descent. This is a direct application of the Representer Theorem (scholkopf2001generalized) and is detailed in Proposition˜1.
@@
Given $f_{t}$}, we may predict the expected feedback for a point $\bm{x}$} as $s({f_{t}}({\bm{x}}))$. Centered around this prediction, we construct confidence sets of the form $s({f_{t}}({\bm{x}}))\pm\beta_{t}(\delta)\sigma_{t}({\bm{x}})]$, and show their uniform anytime validity. The width of the sets are characterized by $\sigma_{t}({\bm{x}})$ defined as
@@
where $\kappa=\sup_{a\leq B}1/\dot{s}(a)$, with $\dot{s}(a)=s(a)(1-s(a))$ denoting the derivative of the sigmoid function, and $K_{t}\in{\mathbb{R}}^{t\times t}$ is the kernel matrix satisfying $K_{t}]_{i,j}=k({\bm{x}}_{i},{\bm{x}}_{j})$. Our first main result shows that for a careful choice of $\beta_{t}(\delta)$, these sets contain $s(f({\bm{x}}))$ simultaneously for all ${\bm{x}}\in{\mathcal{X}}$ and $t\geq 1$ with probability greater than $1-\delta$.
@@
Assume $f\in{\mathcal{H}}_{k}$ and $\left\lVert f\right\rVert_{k}\leq B$. Let $0<\delta<1$ and set
@@
Function-valued confidence sets around the kernelized ridge estimator are analyzed and used extensively to design bandit algorithms with noisy feedback on the true reward values (valko2013finite; srinivas2009gaussian; chowdhury2017kernelized; whitehouse2023improved). However, under noisy logistic feedback, this literature falls short as the proposed confidence sets are no longer valid for the kernelized logistic estimator $f_{t}$}. One could still estimate $f$ using a kernelized ridge estimator and benefit from this line of work. However, as empirically demonstrated in Figure˜1(a), this will not be a sample-efficient approach.
Proof Sketch. When minimizing the kernelized logistic loss, we do not have a closed-form solution for $f_{t}$}, and can only formulate it using the fact that the gradient of the loss evaluated at $f_{t}$} is the null operator, i.e., $\nabla{\mathcal{L}}({f_{t}};H_{t}):{\mathcal{H}}\rightarrow{\mathcal{H}}=\bm{0}$. The key idea of our proof is to construct confidence intervals as $\mathcal{H}$}-valued ellipsoids in the gradient space and show that the gradient operator evaluated at $f$ belongs to it with high probability (c.f. Lemma˜8). We then translate this back into intervals around point estimates $s({f_{t}}({\bm{x}}))$ uniformly for all points ${\bm{x}}\in{\mathcal{X}}$. The complete proof is deferred to Appendix˜A, and builds on the results of faury2020improved and whitehouse2023improved.
Logistic Bandits. Such confidence sets are an integral tool for action selection under uncertainty, and bandit algorithms often rely on them to balance exploration against exploitation. To demonstrate how Theorem˜2 may be used for bandit optimization with direct logistic feedback, we consider LGP-UCB, the kernelized Logistic GP-UCB algorithm. Presented in Algorithm˜2, it extends the optimistic algorithm of faury2020improved from the linear to the kernelized setting, by using the confidence bound of Theorem˜2 to calculate an optimistic estimate of the reward. We proceed to show that LGP-UCB attains a sublinear logistic regret, which is commonly defined as
@@
Let $\delta\in(0,1]$ and choose the exploration coefficients $\beta_{t}(\delta)$ as described in Theorem˜2 for all $t\geq 0$. Then LGP-UCB satisfies the anytime cumulative regret guarantee of